OpenStudy (anonymous):

Please help me on this. Include an explanation and numbers and the answer please. :( I am in the 3rd grade and I dont understand this :( A cookie recipe calls for 2/3 cup of brown sugar. Sarah is making 1/4 of the recipe. How much brown sugar will she need?

OpenStudy (anonymous):

Help. :(

OpenStudy (anonymous):

one fourth of \(\frac{2}{3}\) means \[\frac{1}{4}\times \frac{2}{3}\]

OpenStudy (anonymous):

Why do I multiply?

OpenStudy (anonymous):

one fourth OF something means multiply

OpenStudy (anonymous):

OH! Thanks ill remember that! :D

OpenStudy (anonymous):

The simplified version of the answer wold be 1/6 right?